5×5-5+5÷5=????if you are a genius answer it correctly ok.

Grade:11

9 Answers

Md Saif Ali
6 years ago
we will use BODMAS RULE here.firstly we will divide then multiply then substract and then add .accordingly we will get the answer 21
Riya
6 years ago
Correct answer is 21 by applying BODMASS ruleSo firstly divide,multiply,addition and then subtraction u will get ur correct answer
amisha
6 years ago
in this we will use bodmas
Bracket
Of
Division
Multiplication
Addition
Subtraction
so
5*5-5+5/5
5*5-5+1
25-5+1
26-5
21
hence 21 is the answer
simple....
Sarthak Agarwal
5 years ago
5*5-5+5/55*5-5+125-5+120+121Here we have use rule thus bracket of division multiplication addition and subtraction
ASh
5 years ago
 
in this we will use bodmas
Bracket
Of
Division
Multiplication
Addition
Subtraction
so
5*5-5+5/5
5*5-5+1
25-5+1
26-5
21
hence 21 is the answer
simple...
Sisir
5 years ago
Use BODMAS rule and get finally the result 21.
B= Brackets
O=Order
D=Division
M=Multiplication
A=Addition
S=Subtraction.
Kshitiz Agrawal
5 years ago
So simply by applying BODMAS rule which says Bracket of Division Multiplication Addition and Substraction.
First of all solve the divide symbol i.e. 5/5=1.
Then write the equation as 5*5-5+1
Now solve the multiply sign which gives us 25-5+1. And here you get the answer as 21
Veena
5 years ago
Here we can use BODMAS rule. BODMAS rule explains that while we solve a numerical problem of different sign , first we have to open the bracket if any, the 2nd thing we have to check whether there is any 'OF' , eg: 5 percent OF 100, that is 100×5/100 which is equal to 5. Next part that is the thing we have to use in our problem, ie,DMAS, which means Division, Multiplication, Addition, then Subtraction, So in our question 5*5-5+5/5, first we have to done Division, so it will become
5*5-5+1,as 5/5=1
Then we have to do the multiplication, 25-5+1, as 5*5=25
Then addition have to be done, 25-4, as -5+1=4, here 2 numbers having different sign  , so we have to subtract the small number from the large number and put the sign of the large number, 
Then the last one is subtraction and we get the answer 21.
Guri Bhangu
5 years ago
The right answer according to bodmas rule is21
5*5-5+5/5
So
Firstly divide and we have 
5*5-5+1
Then multiply 
25-5+1
So add 
26-5
In last subtract 
21 that is the answer 
So simple... ... .